Trace the curve y^2 = (x +1) (x −1)^2 by showing all the properties you use to trace it.
1
Expert's answer
2019-08-18T12:57:24-0400

Trace the curve "y^{2}=(x+1)(x-1)^{2}" by showing all the properties you use to trace it.

Solution:

Step 1: Domain

"D(y): (x+1)(x-1)^2\geqslant0"

"(x+1)\geqslant0"

"x\geqslant-1"

"D(y)=[-1;+\infin)"

Step 2: Range

"E(y)=\R"

Step 3: Symmetry

"(-y)^{2}=(x+1)(x-1)^{2}=y^{2}"

The curve is symmetric about x-axis

Step 4: Asymptotes

It has no asymptotes

Step 5: Intersection points with axes:

x-axis: "y=0, x=\pm1"

y-axis: "x=0, y=\pm1"

Step 6: Monotonicity

"y=\pm(x-1)\sqrt{x+1}"

"y'=\pm(\sqrt{x+1}+\frac{x-1}{2\sqrt{x+1}})"

"y'=\pm\frac{2x+2+x-1}{2\sqrt{x+1}}"

"y'=\pm\frac{3x+1}{2\sqrt{x+1}}"

Let "y'=0"

"\pm\frac{3x+1}{2\sqrt{x+1}}=0"

If "x=-1" then "y'=\infin" or "y'" does not exists

"\pm(3x+1)=0"

"x=-\frac{1}{3}"

"y^{2}(-\frac{1}{3})=\frac{2}{3}\cdot\frac{16}{9}=\frac{32}{27}"

"y=\pm\frac{\sqrt{32}}{\sqrt{27}}\approx\pm1.09"

Let "y<0"



So that"(-\frac{1}{3};-\frac{\sqrt{32}}{\sqrt{27}})" is minimum

Let "y>0"


So that "(-\frac{1}{3};\frac{\sqrt{32}}{\sqrt{27}})" is maximum

Step 7: Curve